baconway:

If I need $5,000 within the next 9 months how much would I have to make per work day and how many hours if I only worked 4 days a week and I make 14 an hour

axie:

357 hours for 5k You would need to make 425 dollars a shift to make that much

surjithayer:

money have to be earned per month=5000 /9\approx $ 555.56 number of working days per month=4*4=16 money earned per day=555.56/16\approx $ 34.7 3 number of hors per day=34.73/14 \approx=2.48 1 so he has 2.5 hrs per day earning per day=14*2.5=$35 earning per week=4*35=$140 earning per month=4*140=$560 total earnings=9*560=$ 5040

baconway:

@surjithayer wrote:
money have to be earned per month=5000 /9\approx $ 555.56 number of working days per month=4*4=16 money earned per day=555.56/16\approx $ 34.7 3 number of hors per day=34.73/14 \approx=2.48 1 so he has 2.5 hrs per day earning per day=14*2.5=$35 earning per week=4*35=$140 earning per month=4*140=$560 total earnings=9*560=$ 5040
Thanks that was honestly really helpful and exactly what I was wanting

baconway:

@axie wrote:
357 hours for 5k You would need to make 425 dollars a shift to make that much
Thank you